11-6 Coronado Company purchased equipment for $248,240 on October 1, 2017. It is estimated that the...

11-6

Coronado Company purchased equipment for $248,240 on October 1, 2017. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $13,920. Estimated production is 40,400 units and estimated working hours are 20,000. During 2017, Coronado uses the equipment for 530 hours and the equipment produces 1,100 units.

Compute depreciation expense under each of the following methods. Coronado is on a calendar-year basis ending December 31. (Round rate per hour and rate per unit to 2 decimal places, e.g. 5.35 and final answers to 0 decimal places, e.g. 45,892.)

(a) Straight-line method for 2017 $

(b) Activity method (units of output) for 2017 $

(c) Activity method (working hours) for 2017 $

(d) Sum-of-the-years'-digits method for 2019 $

(e) Double-declining-balance method for 2018 $
Answer #1

A. Straight line method.

>> Depreciation per annum = ( $ 248,240 - $ 13,920 ) / 8 = $ 29,290

>> Depreciation for 2017 = $ 29,290 * ( 3 /12 ) = $ 7,323

________________________________________________________________________________________

B. Units of output.

>> Depreciation per unit = ( $ 248,240 - $ 13,920 ) / 40,400 = $ 5.8

>> Depreciation for 2017 = 1,100 * $ 5.8 = $ 6,380.

__________________________________________________________________________________________

C. Working ours.

>> Depreciation per hours = ( $ 248,240 - $ 13,920 ) / 20,000 = $11.716

>> Depreciation for 2017 = 530 * $ 11.716 = $ 6,210

______________________________________________________________________________________________

D. Sum of digits method

>> Sum of years = 8 ( 8 +1 ) / 2 = 36.

>> Depreciation for 2019 =( $ 248,240 - $ 13,920 ) * ( 6.75 / 36 ) = $ 43,935

___________________________________________________________________________________________

E . Double declining balance

>> Depreciation rate = 200 / 8 = 25 %.

>> Depreciation for 2017 = $ 248,240 * 25 % * ( 3 /12 ) = $ 15,515

>> Depreciation for 2018 = ( $ 248,240 - $ 15,515 )* 25 % = $ 58,181